OpenStudy (anonymous):

Amy is getting in shape for the track season. Her coach suggests running a certain number of kilometers each week. Amy runs laps around a lake near her house to reach this goal. Here's the formula for this scenario: n = \[\frac{ k }{ d }\] k = total number of kilometers d = distance around the lake n = number of laps Amy runs around the lake Rewrite the formula to determine the distance around the lake.

OpenStudy (hihi67):

d = k/n

OpenStudy (hihi67):

@Skiller8860 Do you get it?

OpenStudy (hihi67):

Look: n = k/d Times d on both sides. n*d = k Divide by n on both sides. d = k/n

OpenStudy (anonymous):

Not really getting it. And sorry was away....

OpenStudy (hihi67):

It's just an equation.

OpenStudy (hihi67):

What grade are you in?

OpenStudy (anonymous):

I understand that, however I do not get what I plug into the variables.

OpenStudy (anonymous):

I'm in 10th. Doesn't matter.

OpenStudy (hihi67):

Why not? Lets say k is 10 and d is 2.

OpenStudy (hihi67):

So k / d would be 5. So n is 5.

OpenStudy (anonymous):

I understand, but does it want me to rewrite it in a different format, or does it want me to figure out what to plug into the variables and solve for it?

OpenStudy (hihi67):

Then k / n should be d, so 10 / 5 should be 2, and it is...

OpenStudy (hihi67):

Noooo, just rewrite it.

OpenStudy (anonymous):

Oh .. Hah'. Thanks xD.

OpenStudy (hihi67):

No problem. :D